OpenStudy (anonymous):

log3 x + log3(x-24)=4

OpenStudy (anonymous):

\[a. \left\{ 27 \right\} b. \left\{ 53 \right\} c.\left\{ -3,27 \right\}d.\emptyset\] choose what best completes the statement